OpenStudy (anonymous):

3. The rectangular prism has a volume of 105 m3. This prism is enlarged using a scale factor of 4, then What is the volume of the enlarged prism? A. 6720 m3 B. 1680 m3 C. 840 m3 D. 420 m3

OpenStudy (anonymous):

think of the equation of the volume of a rectangle.

OpenStudy (anonymous):

V = L * W * H

OpenStudy (anonymous):

if you enlarge it by a factor of 4, what happens to its L, W, and H

OpenStudy (nurali):

volume = lbh = 105 --------(1) when u enlarge it wid a scalefactor of 4, you're multiplying each dimension wid 3. so, new volume = (4l)(4b)(4h) =4^3 lbh =64 x 105 = ?

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

is the scale factor a linear scale factor or volume scale factor?

OpenStudy (anonymous):

if it is a linear one the use the formula v.s.f=l.s.f^3 then multiply the v.s.f with the volume to get the new volume.